### v1

# Coherent Information Phase Transition in a Noisy Quantum Circuit

Abstract
Coherent information quantifies the transmittable quantum information through a channel and is directly linked to the channel’s quantum capacity. In the context of dynamical purification transitions, scrambling dynamics sustain extensive and positive coherent information at low measurement rates, but noises can suppress it to zero or negative values. Here we introduce quantum-enhanced operations into a noisy monitored quantum circuit. This circuit, viewed as a quantum channel, undergoes a phase transition in coherent information from a recoverable phase with positive values to an irrecoverable phase with negative values. This transition is modulated by the relative frequency of noise and quantum-enhanced operations. The existence of a recoverable phase implies that quantum-enhanced operations can facilitate reliable quantum information transmission in the presence of diverse noises. Remarkably, we propose a resource-efficient protocol to characterize this phase transition, effectively avoiding post-selection by utilizing every run of the quantum simulation. This approach bridges the gap between theoretical insights and practical implementation, making the phase transition feasible to demonstrate on realistic noisy intermediate-scale quantum devices.
Quantum information is fundamentally represented by quantum entanglement, which serves as a critical resource in both quantum computation and quantum communication [1]. In realistic experimental settings, quantum systems are typically modeled as open systems, where entanglement between different parts of the system is frequently transferred to the surrounding environment due to decoherence [2, 3]. This transfer, coupled with the principle of entanglement monogamy, results in an irreversible loss of information to the environment, thereby diminishing the potential quantum advantages [4]. To quantify the extent of information loss during quantum channel transmission, coherent information stands as a key metric, closely linked to the quantum channel capacity [5, 6, 7, 8, 9]. Positive coherent information indicates the successful transmission of finite quantum information through a channel, whereas zero or negative values suggest that no quantum information is being transmitted. As such, the development of methods to maintain positive coherent information in the presence of noise is a pivotal area of research in the quest for fault-tolerant quantum computation [10].
One approach to achieving positive coherent information involves encoding quantum information within an enlarged Hilbert space, transmitting it through a noisy quantum channel, and subsequently decoding it--this encapsulates the essence of quantum error correction [11, 12, 13, 14, 15]. By utilizing only a subspace of the total Hilbert space to represent logical information, errors can be effectively detected and corrected. An alternative method involves encoding information in a highly non-local manner within the same Hilbert space, leveraging quantum scrambling [16, 17, 18]. In the context of measurement-induced phase transitions (MIPT), it has been shown that a low rate of local measurements is insufficient to extract significant information when competing with scrambling dynamics generated by random unitary gates [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. This phenomenon is more transparently elucidated by relating MIPT to a dynamical purification transition [39], characterized by circuit-averaged coherent information, where the input state is a completely mixed state. In such scenarios, coherent information is extensive in system size and remains positive in the mixed phase, while it approaches zero in the pure phase. However, it is important to note that scrambling alone is insufficient to protect information from other prevalent sources of noise and may even aggravate the suppression of entanglement [40, 41, 42, 43, 44, 45]. As a result, negative coherent information is expected in noisy quantum circuits involving random unitary gates and measurements.
In this Letter, we explore the integration of quantum-enhanced (QE) operations into quantum circuits, unveiling a phase transition in coherent information that is governed by the relative frequency of various noises and QE operations. Coherent information remains positive in the recoverable phase but becomes negative in the irrecoverable phase. QE operations involve utilizing a quantum probe to extract information from the system, which in our framework can be conceptualized as a dynamic expansion of the Hilbert space through the introduction of ancilla qubits during the circuit evolution [46, 47, 48, 49, 50, 51, 52]. We further demonstrate that this phase transition can be efficiently probed in experiments by generalizing the cross entropy benchmark for MIPT [53], thereby circumventing the challenges associated with post-selection. This finding underscores that the phase transition remains experimentally tractable even in the thermodynamic limit, where phases of matter are well-defined [54].
Figure 1: Circuit structure and phase diagram. (a) Circuit structure. Orange, green, yellow, and blue rectangles represent unitary gates, measurements, noise, and QE operations, respectively. The initial state is entangled with reference qubits $R$. We primarily consider QE operations where $U_{SA}=\text{SWAP}$. (b) Phase diagram. We focus on the first-order transition from recoverable to irrecoverable phase, tuned by the relative frequency of noise and QE operations.
Circuit model. We consider a quantum circuit structure composed of four types of operations, as depicted in Fig. 1(a). Random unitary gates are applied in a brick-wall pattern. Between each pair of unitary layers, each qubit has a probability $p$ of being projectively measured along the $z$-axis. The projective measurement can be modeled by first applying a CNOT gate between the system qubit and an environment qubit, followed by another CNOT gate with an ancilla qubit. Tracing out the environment qubit leaves the classical measurement outcome in the ancilla qubit. After the measurement, each qubit has a probability $q_{n}$ of undergoing a certain noise channel, such as depolarizing, resetting, or dephasing, and a probability $q_{e}$ of undergoing a QE operation. We denote $q_{t}=q_{n}+q_{e}$ and $q=q_{n}/q_{t}$. The QE operation involves introducing ancilla qubits and then applying a unitary gate $U_{SA}$ to the system qubit and these ancilla qubits. These ancilla qubits are then isolated, remaining untouched until the end of the circuit. Given that each ancilla qubit is used only once, we assume no noise affects them during the evolution, which is a reasonable assumption considering they can be well-isolated from other qubits, thereby maintaining a coherence time much longer than that of the system qubits. In Refs. [50, 51, 52], the case where noise and QE operations are symmetric was studied. Here, we consider the relative frequency of noise and QE operations, $q$, is no longer constrained to be $0.5$.
The input state is entangled with a reference system $R$, and the entire quantum circuit can be viewed as a quantum channel from $R$ to $SA$. Without loss of generality, we consider the scenario where each system qubit is entangled in a Bell pair with a corresponding reference qubit. It is important to note that $A$ includes both ancilla qubits from QE operations that store quantum information and those from measurements that only contain classical information. Coherent information, a key quantity for assessing quantum channel capacity, is computed as:
$I_{C}(R>SA)=\mathcal{S}(SA)-\mathcal{S}(SRA),$ (1)
where $\mathcal{S}(X)$ denoted the entanglement entropy of the subsystem $X$. Our primary focus is to investigate whether a phase transition in $I_{C}(R>SA)$ occurs as $q$ is varied.
Analytical analysis. When random unitary gates are drawn from the Haar measure, the coherent information can be mapped onto the free energy difference of a classical statistical mechanics model under different boundary conditions [55, 56, 57, 58]. Specifically, we have
$I_{C}(R>SA)$ $=$ $\lim_{n\to 1}(\mathcal{S}^{(n)}(SA)-\mathcal{S}^{(n)}(SRA))$
$=$ $\lim_{n\to 1}\lim_{k\to 0}\frac{1}{(1-n)k}\text{log}\left(\frac{\mathcal{Z}^{(n,k)}_{SA}}{\mathcal{Z}^{(n,k)}_{SRA}}\right)$
$=$ $\lim_{n\to 1}\lim_{k\to 0}\frac{1}{(n-1)k}\left(\mathcal{F}^{(n,k)}_{SA}-\mathcal{F}^{(n,k)}_{SRA}\right),$
where
$\mathcal{Z}_{XA}^{(n,k)}$ $=$ $\sum_{m}\text{Tr}\left(\mathbb{C}^{X}\Lambda_{A}^{(n,k)}\right),\quad\mathcal{F}_{XA}^{(n,k)}=-\text{log}\mathcal{Z}_{XA}^{(n,k)},$
$\Lambda_{A}^{(n,k)}$ $=$ $\text{Tr}_{A\cup E}\left[\left(\left|\Psi_{m}\right\rangle\left\langle\Psi_{m}\right|\right)^{\otimes Q}\mathbb{C}^{A}\right].$ (3)
Here, the degrees of freedom of this statistical mechanics model are spins that take values in the permutation group $\mathbb{S}(Q)$ with $Q=nk+1$, and are denoted as $\left|\sigma\right\rangle$. $X=S,SR$. $\mathbb{C}$ corresponds to a particular group member in $\mathbb{S}(Q)$, representing to a particular spin orientation. Thus, $\Lambda_{A}^{(n,k)}$ can be identified as the bulk partition function, while $\mathbb{C}^{S}$ and $\mathbb{C}^{SR}$ correspond to two different boundary conditions. Notice that $S$ and $R$ represents the top and bottom boundary, respectively.
This statistical mechanics model is a ferromagnetically coupled spin model with random fields oriented in two different directions on a honeycomb lattice [52]. The vertical bond contributes a weight of $\text{Wg}^{Q}(\sigma^{-1}\tau)$, where $\text{Wg}^{Q}(g)$ is the Weingarten function, acting as a ferromagnetic coupling that assigns greater weights to configurations where $\sigma=\tau$. The non-vertical bonds are determined by the specific choice of QE operations and noise. For instance, if we select $U_{SA}=\text{SWAP}$ and consider depolarizing noise, the bond contributes a weight:
$W^{\text{depo}}$ $(\sigma,\tau)=(1-p)(1-q_{t})\left\langle\sigma|\tau\right\rangle+dp(1-qq_{t})+qq_{t}\times$
$[(1-p)\left\langle\sigma|\mathbb{I}\right\rangle+dp]\left\langle\mathbb{I}|\tau\right\rangle+(1-p)(1-q)q_{t}\left\langle\sigma|\mathbb{C}\right\rangle.$
Here, $\left\langle\sigma|\tau\right\rangle=d^{|\sigma\tau^{-1}|}$, where $d$ is the local Hilbert space dimension and where $|\sigma\tau^{-1}|$ denotes the number of cycles in permutation $\sigma\tau^{-1}$. $\mathbb{I}$ corresponds to the identity element in $\mathbb{S}(Q)$. Thus, $q$ determines the net field direction. For small $q$, the net field aligns with $\mathbb{C}$, whereas for large $q$, it aligns with $\mathbb{I}$. When combined with the results in Ref. [52], the full phase diagram is depicted in Fig. 1(b). By identifying $p$ as the temperature and considering noise and QE operations as two competing external fields, the phase diagram is reminiscent of that of a simple 2D Ising model. Consequently, a first-order phase transition is expected when tuning $q$.
Furthermore, the coherent information can directly detect this first-order phase transition. For small $q$, bulk spins align with $\mathbb{C}$, leading to a domain wall at the bottom boundary when calculating $\mathcal{F}^{(n,k)}_{SA}$, with no additional energy cost in $\mathcal{F}^{(n,k)}_{SAR}$. This results in a positive $I_{C}$ and is extensive in system size. Conversely, for large $q$, bulk spins align with $\mathbb{I}$. In this scenario, a domain wall at the top boundary always exists, while an extra domain wall appears in calculating $\mathcal{F}^{(n,k)}_{SAR}$ at the bottom boundary, rendering $I_{C}$ negative. Since positive $I_{C}$ indicates that finite amount of quantum information can be successfully transmitted through the channel, which can be recovered via a decoding algorithm, this first-order phase transition corresponds to a transition from a recoverable to an irrecoverable phase.
Figure 2: Numerical simulation for coherent information phase transition. (a) Resetting noise. (b) Depolarizing noise. (c) Dephasing noise. (d) Depolarizing noise with measurement probablility $p=0.1$. Insets show the data collapse results. Every data point is averaged over $6\times 10^{3}$ realizations. The circuit is evolved for $5L$ time steps.
Numerical results. To explicitly demonstrate the existence of the coherent information phase transition, we perform numerical simulations on a large scale. In these simulations, random unitary gates are selected from the Clifford group to enable the use of the stabilizer formalism [59, 60, 61]. We set $U_{SA}=\text{SWAP}$ and specifically consider three types of noise: resetting, depolarizing, and dephasing. Here we set $q_{t}=0.1$, while results for other parameters can be found in Supplemental Material [62]. We first consider the case with $p=0$, and the results are presented in Fig. 2(a)-2(c). A distinct recoverable phase, characterized by positive coherent information, is identified. Through data collapse, we determine the critical points and critical exponents to be $q_{c}^{\text{reset}}=0.500(1)$, $q_{c}^{\text{depo}}=0.360(1)$ and $q_{c}^{\text{deph}}=0.503(1)$, with $\nu_{c}^{\text{reset}}=0.90(5)$, $\nu_{c}^{\text{depo}}=0.96(5)$ and $\nu_{c}^{\text{deph}}=0.95(7)$, respectively. The recoverable phase region is smallest for depolarizing noise and largest for dephasing noise, consistent with the relative strengths of the corresponding random fields [62]. When measurements are incorporated, the phase transition persists, as demonstrated in Fig. 2(d) for depolarizing noise, albeit at a different critical point $q_{c}=0.375(1)$ with $\nu_{c}=1.03(8)$. Additional numerical details and results are provided in [62].
Figure 3: Numerical evidence of critical slowing down. We consider resetting noise with $p=0$ as an example. (a) Temporal evolution of $I_{C}$ for varying $q$. We take system size to be $L=256$. (b) Convergence time vs. $q$ for multiple system sizes. (c), (d) analogous to (a), (b), but in absence of random unitary gates.
Another interesting phenomenon we observed is critical slowing down [63, 64, 65, 66]. As illustrated in Fig. 3(a), the convergence of the coherent information slows significantly near $q=0.5$ for resetting noise. By defining the convergence time $t_{c}$ as the time when $\left|I_{C}(t)-I_{C}(t_{\text{final}})\right|$ first falls below $0.05$, we plot $t_{c}$ as a function of $q$ for different system sizes in Fig. 3(b). The results indicate that $t_{c}$ diverges with system size as $q$ approaches the critical point, suggesting an infinite convergence time in the thermodynamic limit. In contrast, in the absence of unitary gates, while the coherent information still transitions from positive to negative precisely at $q_{c}=0.5$ (corresponding to the scenario where half of the system qubits are discarded), critical slowing down was not observed, as shown in Fig. 3(c) and 3(d). This observation underscores the indispensability of unitary gates for the manifestation of a phase transition.
Efficient protocol. Measuring coherent information for a specific circuit and trajectory requires repeated preparation of the same final state. However, in realistic experiments, noise locations are usually uncontrollable, making it difficult to reliably reproduce the same circuit. This issue is compounded by the post-selection problem when measurements are involved, due to Born’s rule [67, 68, 54]. Motivated by previous work on constructing post-selection-free probes that reflect the correlation between a quantum device and classical simulations [69, 53, 70], we propose a resource-efficient protocol to observe the coherent information phase transition. In the quantum device, an initial state $\rho$ is prepared and encoded by the application of random unitary gates alone. This encoded state then undergoes the circuit, with ancillary qubits introduced during the evolution. The unnormalized final state on $A$ is denoted as $\widetilde{\rho_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\rho_{A}^{m}})=p_{\rho}^{m}$, where $p_{\rho}^{m}$ represents the probability of the particular trajectory $m$. For brevity, we omit the explicit dependence on the circuit realization $\mathcal{C}$, which includes choices of random unitary gates and locations of noise, QE operations and measurements. Leveraging the knowledge of the measurement outcomes and the circuit realization, we execute a classical simulation of the same circuit, but with the input state being the encoded state of a distinct state $\sigma$. Assuming the unitary gates are Clifford gates and $\sigma$ is a stabilizer state, the classical simulation becomes computationally efficient. We then also obtain $\widetilde{\sigma_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\sigma_{A}^{m}})=p_{\sigma}^{m}$. It is worth noting that the initial state $\rho$ can be chosen beyond stabilizer states, making the evolution intractable on a classical computer.
Figure 4: Efficient protocol. (a) Schematic workflow of the protocol. (b) Take depolarizing noise as an example. We choose $\rho=(\left|+\right\rangle\left\langle+\right\rangle)^{\otimes L/2}\otimes(\left|0\right\rangle\left\langle 0\right\rangle)^{\otimes L/2}$ and $\sigma=(\left|0\right\rangle\left\langle 0\right\rangle)^{\otimes L}$. Every data point is averaged over $3\times 10^{3}$ realizations. The inset shows the data collapse result. (c) Analogous to (b), with measurement probability $p=0.1$.
The quantity that serves as a proxy for the phase transition is constructed as follows:
$\chi$ $=$ $\mathbb{E}_{\mathcal{C}}\frac{\sum_{m}\text{Tr}\left(\widetilde{\rho_{A}^{m}}\widetilde{\sigma_{A}^{m}}\right)}{\sum_{m}\text{Tr}\left(\widetilde{\sigma_{A}^{m}}^{2}\right)}$ (4)
$=$ $\mathbb{E}_{\mathcal{C}}\sum_{m}p_{\rho}^{m}\frac{p_{\sigma}^{m}}{2^{-N_{\text{rand}}^{\sigma}}}\text{Tr}\left(\rho_{A}^{m}\prod_{k=1}^{l_{\sigma}}\frac{1+g_{k}}{2}\right).$
Here, $\mathbb{E}_{\mathcal{C}}$ represents averaging over different circuit realizations, $N_{\text{rand}}^{\sigma}$ is the number of measurements with random outcomes during the classical simulation, $l_{\sigma}$ represents the number of generators for the normalized stabilizer state $\sigma_{A}^{m}$ and $g$ are the corresponding generators. The middle term $p_{\sigma}^{m}/2^{-N_{\text{rand}}^{\sigma}}$ is either 00 or $1$ due to the property of stabilizer formalism, while the final term corresponds to the probability of measuring all the generators with outcomes equal to $1$. For each execution of the quantum circuit, we first compute $p_{\sigma}^{m}/2^{-N_{\text{rand}}^{\sigma}}$ by classical simulation. If the result is $1$, we proceed to perform projective measurements on the ancilla qubits according to the stabilizer generators. If all measurement outcomes are $1$, we denote this as a successful event. The quantity $\chi$ is then estimated as the proportion of successful events among all circuit executions. The whole procedure is outlined in Fig. 4(a). Remarkably, this approach allows us to fully leverage each circuit realization and trajectory, circumventing the need for experimental control over noise locations or post-selection on measurement outcomes.
Physically, this probe reflects the distinguishability of different initial states by leveraging access to both the measurement outcomes and the final state of the ancilla qubits. In the recoverable phase, the initial information is preserved, and we therefore expect $\chi=0$, corresponding to perfect distinguishability. Conversely, in the irrecoverable phase, the initial information becomes erased, causing $\chi$ to approach 1. A detailed analytical analysis of this protocol is provided in [62]. To demonstrate the efficacy of this probe, we present numerical results in Fig. 4(b) and Fig. 4(c) for the case of depolarizing noise with $p=0$ and $p=0.1$, respectively. The critical points are determined to be $q_{c}=0.362(2)$ and $q_{c}=0.378(2)$, which are in agreement with the exact phase transition points.
Discussions. The coherent information is related to the single-shot quantum channel capacity by the relation $\mathcal{Q}_{1}=\max_{\rho_{SR}}I_{C}(R>SA)$ [9, 2]. Here we consider $\rho_{SR}$ to be a direct product of Bell pairs between $S$ and $R$. Through analytical mapping, it is evident that this choice maximizes the coherent information within the recoverable phase, minimizing the free energy $\mathcal{F}^{(n,k)}_{SAR}$. However, in the irrecoverable phase, selecting $\rho_{SR}$ as a trivial product state would yield zero coherent information, thereby maximizing the coherent information. This results in a phase transition where the single-shot quantum channel capacity shifts from a finite positive value to zero. Furthermore, the true quantum channel capacity $\mathcal{Q}$, which allows for multiple uses of the channel, becomes additive and equal to $\mathcal{Q}_{1}$ only if the channel is degradable [71, 72, 73, 74, 75, 76, 77]. Although measurement and QE operations are all degradable, general noises often lacks this property, making the calculation of $\mathcal{Q}$ more challenging. Thus, it is intriguing to explore whether a phase transition in $\mathcal{Q}$ can be identified within this framework.
[…]
Acknowledgements.
Acknowledgment. We are deeply indebted to Xiao-Liang Qi for many valuable discussions. This work is supported by the National Key Research Program of China under Grant No. 2019YFA0308404, the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600 and No. 2019SHZDZX01.

### v2

@@
 We focus on the first-order transition from recoverable to irrecoverable phase, tuned by the relative frequency of noise and QE operations.
+This phase transition manifests irrespective of the measurement rate $p$.
 Circuit model.
@@
 The vertical bond contributes a weight of $\text{Wg}^{Q}(\sigma^{-1}\tau)$, where $\text{Wg}^{Q}(g)$ is the Weingarten function, acting as a ferromagnetic coupling that assigns greater weights to configurations where $\sigma=\tau$.
-The non-vertical bonds are determined by the specific choice of QE operations and noise.
-For instance, if we select $U_{SA}=\text{SWAP}$ and consider depolarizing noise, the bond contributes a weight:
-$W^{\text{depo}}$ $(\sigma,\tau)=(1-p)(1-q_{t})\left\langle\sigma|\tau\right\rangle+dp(1-qq_{t})+qq_{t}\times$
-$[(1-p)\left\langle\sigma|\mathbb{I}\right\rangle+dp]\left\langle\mathbb{I}|\tau\right\rangle+(1-p)(1-q)q_{t}\left\langle\sigma|\mathbb{C}\right\rangle.$
-Here, $\left\langle\sigma|\tau\right\rangle=d^{|\sigma\tau^{-1}|}$, where $d$ is the local Hilbert space dimension and where $|\sigma\tau^{-1}|$ denotes the number of cycles in permutation $\sigma\tau^{-1}$.
-$\mathbb{I}$ corresponds to the identity element in $\mathbb{S}(Q)$.
-Thus, $q$ determines the net field direction.
+The non-vertical bonds are determined by measurements, QE operations and noise.
+While their exact contributions depend on the specific QE operations and noise model, QE operations generally behave as external fields applied at random locations, aligning in the direction $\mathbb{I}$, corresponding to the identity element in $\mathbb{S}(Q)$.
+Conversely, noise acts as an external field oriented in the direction $\mathbb{C}$.
+Concrete expressions for the bond weights under various noises are provided in Ref.
+[59].
+The parameter $q$ governs the relative frequency of QE operations and noise, thereby tuning the net field direction.
 For small $q$, the net field aligns with $\mathbb{C}$, whereas for large $q$, it aligns with $\mathbb{I}$.
@@
 In contrast, in the absence of unitary gates, while the coherent information still transitions from positive to negative precisely at $q_{c}=0.5$ (corresponding to the scenario where half of the system qubits are discarded), critical slowing down was not observed, as shown in Fig.
-3(c) and 3(d).
+3(c) and Fig.
+3(d).
 This observation underscores the indispensability of unitary gates for the manifestation of a phase transition.
@@
 This issue is compounded by the post-selection problem when measurements are involved, due to Born’s rule [67, 68, 54].
-Motivated by previous work on constructing post-selection-free probes that reflect the correlation between a quantum device and classical simulations [69, 53, 70], we propose a resource-efficient protocol to observe the coherent information phase transition.
-In the quantum device, an initial state $\rho$ is prepared and encoded by the application of random unitary gates alone.
-This encoded state then undergoes the circuit, with ancillary qubits introduced during the evolution.
+Motivated by previous work on constructing post-selection-free probes that reflect the correlation between a quantum device and classical simulations [69, 53, 70, 71, 72], we propose a resource-efficient protocol to observe the coherent information phase transition.
+In the quantum device, an initial state $\rho$ is prepared and encoded using random unitary gates alone.
+This encoded state then undergoes the circuit, incorporating ancilla qubits during the process.
 The unnormalized final state on $A$ is denoted as $\widetilde{\rho_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\rho_{A}^{m}})=p_{\rho}^{m}$, where $p_{\rho}^{m}$ represents the probability of the particular trajectory $m$.
 For brevity, we omit the explicit dependence on the circuit realization $\mathcal{C}$, which includes choices of random unitary gates and locations of noise, QE operations and measurements.
-Leveraging the knowledge of the measurement outcomes and the circuit realization, we execute a classical simulation of the same circuit, but with the input state being the encoded state of a distinct state $\sigma$.
+Leveraging the knowledge of the measurement outcomes and the circuit realization, we simulate the circuit classically with an encoded version of a distinct input state $\sigma$.
 Assuming the unitary gates are Clifford gates and $\sigma$ is a stabilizer state, the classical simulation becomes computationally efficient.
-We then also obtain $\widetilde{\sigma_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\sigma_{A}^{m}})=p_{\sigma}^{m}$.
-It is worth noting that the initial state $\rho$ can be chosen beyond stabilizer states, making the evolution intractable on a classical computer.
+We then also obtain $\widetilde{\sigma_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\sigma_{A}^{m}})=p_{\sigma}^{m}$ [73].
 Figure 4: Efficient protocol.
 (a) Schematic workflow of the protocol.
+An important additional ingredient, compared to the approach in [53], is the incorporation of projective measurements on the ancilla qubits.
 (b) Take depolarizing noise as an example.
@@
 Remarkably, this approach allows us to fully leverage each circuit realization and trajectory, circumventing the need for experimental control over noise locations or post-selection on measurement outcomes.
+The proxy $\chi$ resembles the “quantum” cross entropy introduced in Ref.
+[53], with the notable distinction that ancilla qubits are introduced dynamically in this context.
+Furthermore, while this proxy does not reveal MIPT, it proves to be a robust indicator of the coherent information transition.
+As demonstrated in Supplemental Material [59], $\chi$, akin to the coherent information, is directly related to the free energy difference between distinct boundary conditions imposed on the bottom boundary.
+This dependency drives a sharp transition in $\chi$ from 00 to $1$.
 Physically, this probe reflects the distinguishability of different initial states by leveraging access to both the measurement outcomes and the final state of the ancilla qubits.
@@
 Conversely, in the irrecoverable phase, the initial information becomes erased, causing $\chi$ to approach 1.
-A detailed analytical analysis of this protocol is provided in [62].
 To demonstrate the efficacy of this probe, we present numerical results in Fig.
@@
 We are deeply indebted to Xiao-Liang Qi for many valuable discussions.
-This work is supported by the National Key Research Program of China under Grant No.
-2019YFA0308404, the Natural Science Foundation of China through Grants No.
+This work is supported by the Natural Science Foundation of China through Grants No.
 12350404 and No.
@@
 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No.
-23JC1400600 and No.
+23JC1400600, No.
+24LZ1400100 and No.
 2019SHZDZX01.

### v3

@@
 Coherent information quantifies the transmittable quantum information through a channel and is directly linked to the channel’s quantum capacity.
-In the context of dynamical purification transitions, scrambling dynamics sustain extensive and positive coherent information at low measurement rates, but noises can suppress it to zero or negative values.
-Here we introduce quantum-enhanced operations into a noisy monitored quantum circuit.
-This circuit, viewed as a quantum channel, undergoes a phase transition in coherent information from a recoverable phase with positive values to an irrecoverable phase with negative values.
-This transition is modulated by the relative frequency of noise and quantum-enhanced operations.
-The existence of a recoverable phase implies that quantum-enhanced operations can facilitate reliable quantum information transmission in the presence of diverse noises.
-Remarkably, we propose a resource-efficient protocol to characterize this phase transition, effectively avoiding post-selection by utilizing every run of the quantum simulation.
+In a monitored quantum circuit, regarded as a quantum channel, extensive and positive coherent information is sustained at low measurement rates, protected by the scrambling dynamics.
+However, noise suppresses coherent information, driving it to zero or negative values.
+Here, we show that incorporating quantum-enhanced operations facilitates reliable quantum information transmission even in the presence of noise, as evidenced by a phase transition in coherent information from a recoverable phase with positive values to an irrecoverable phase with negative values.
+We provide both analytical understanding and numerical evidence demonstrating this transition, which is modulated by the relative frequencies of noise and quantum-enhanced operations.
+Additionally, we propose a resource-efficient protocol to characterize this phase transition in experiments, effectively avoiding post-selection by utilizing every run of the quantum circuit.
 This approach bridges the gap between theoretical insights and practical implementation, making the phase transition feasible to demonstrate on realistic noisy intermediate-scale quantum devices.
 Quantum information is fundamentally represented by quantum entanglement, which serves as a critical resource in both quantum computation and quantum communication [1].
-In realistic experimental settings, quantum systems are typically modeled as open systems, where entanglement between different parts of the system is frequently transferred to the surrounding environment due to decoherence [2, 3].
-This transfer, coupled with the principle of entanglement monogamy, results in an irreversible loss of information to the environment, thereby diminishing the potential quantum advantages [4].
-To quantify the extent of information loss during quantum channel transmission, coherent information stands as a key metric, closely linked to the quantum channel capacity [5, 6, 7, 8, 9].
+However, the transmission of quantum information is inevitably disrupted by interactions with the surrounding environment, leading to decoherence, which diminishes the potential advantages of quantum systems [2, 3].
+To quantify the extent of information loss during quantum channel transmission, coherent information is a key metric, closely related to the quantum channel capacity [4, 5, 6, 7, 8].
 Positive coherent information indicates the successful transmission of finite quantum information through a channel, whereas zero or negative values suggest that no quantum information is being transmitted.
 As such, the development of methods to maintain positive coherent information in the presence of noise is a pivotal area of research in the quest for fault-tolerant quantum computation [9].
-One approach to achieving positive coherent information involves encoding quantum information within an enlarged Hilbert space, transmitting it through a noisy quantum channel, and subsequently decoding it--this encapsulates the essence of quantum error correction [11, 12, 13, 14, 15].
-By utilizing only a subspace of the total Hilbert space to represent logical information, errors can be effectively detected and corrected.
+One approach to achieving positive coherent information involves encoding quantum information within an enlarged Hilbert space, transmitting it through a noisy quantum channel, and subsequently decoding it-this encapsulates the essence of quantum error correction [10, 11, 12, 13, 14].
 An alternative method involves encoding information in a highly non-local manner within the same Hilbert space, leveraging quantum scrambling [15, 16, 17].
@@
 In such scenarios, coherent information is extensive in system size and remains positive in the mixed phase, while it approaches zero in the pure phase.
-However, it is important to note that scrambling alone is insufficient to protect information from other prevalent sources of noise and may even aggravate the suppression of entanglement [40, 41, 42, 43, 44, 45].
+However, it is important to note that scrambling alone is insufficient to protect information transmission from other prevalent sources of noise and may even aggravate the suppression of entanglement [39, 40, 41, 42, 43, 44, 45, 46].
 As a result, negative coherent information is expected in noisy quantum circuits involving random unitary gates and measurements.
-In this Letter, we explore the integration of quantum-enhanced (QE) operations into quantum circuits, unveiling a phase transition in coherent information that is governed by the relative frequency of various noises and QE operations.
-Coherent information remains positive in the recoverable phase but becomes negative in the irrecoverable phase.
-QE operations involve utilizing a quantum probe to extract information from the system, which in our framework can be conceptualized as a dynamic expansion of the Hilbert space through the introduction of ancilla qubits during the circuit evolution [46, 47, 48, 49, 50, 51, 52].
-We further demonstrate that this phase transition can be efficiently probed in experiments by generalizing the cross entropy benchmark for MIPT [53], thereby circumventing the challenges associated with post-selection.
-This finding underscores that the phase transition remains experimentally tractable even in the thermodynamic limit, where phases of matter are well-defined [54].
+In this Letter, we explore the integration of quantum-enhanced (QE) operations [47, 48, 49, 50, 51, 52, 53] into quantum circuits, unveiling a phase transition in coherent information that is governed by the relative frequency of various noises and QE operations.
+In the recoverable phase, the coherent information remains positive, indicating reliable quantum information transmission even in the presence of noise, while in the irrecoverable phase, it becomes negative.
+Our results thus demonstrate that QE operations are effective not only in storing quantum information [53], but also in transmitting it.
+This transition can be analytically understood by mapping the circuit to a classical statistical mechanics model and is explicitly demonstrated through numerical simulations.
+Moreover, we demonstrate that this phase transition can be efficiently probed in experiments by generalizing the cross-entropy benchmark for MIPT [54], thereby circumventing the challenges associated with post-selection.
+This finding underscores that the phase transition remains experimentally tractable even in the thermodynamic limit, where the phases of matter are well defined [55].
 Figure 1: Circuit structure and phase diagram.
@@
 (b) Phase diagram.
-We focus on the first-order transition from recoverable to irrecoverable phase, tuned by the relative frequency of noise and QE operations.
-This phase transition manifests irrespective of the measurement rate $p$.
-Circuit model.
-We consider a quantum circuit structure composed of four types of operations, as depicted in Fig.
+We focus on the coherent information phase transition from recoverable to irrecoverable phase, tuned by the relative frequency of noise and QE operations.
+This phase transition manifests irrespective of the measurement rate pp.
+Circuit model—We consider a quantum circuit structure composed of four types of operations, as depicted in Fig.
 1(a).
@@
 We denote $q_{t}=q_{n}+q_{e}$ and $q=q_{n}/q_{t}$.
-The QE operation involves introducing ancilla qubits and then applying a unitary gate $U_{SA}$ to the system qubit and these ancilla qubits.
-These ancilla qubits are then isolated, remaining untouched until the end of the circuit.
+QE operations involve utilizing a quantum probe to extract information from the system, which can be conceptualized as a dynamic expansion of the Hilbert space through the introduction of ancilla qubits during the circuit evolution.
+Specifically, a QE operation introduces ancilla qubits and applies a unitary gate to the system qubit and these ancilla qubits, which are then isolated and left untouched until the end of the circuit.
 Given that each ancilla qubit is used only once, we assume no noise affects them during the evolution, which is a reasonable assumption considering they can be well-isolated from other qubits, thereby maintaining a coherence time much longer than that of the system qubits.
-In Refs.
-[50, 51, 52], the case where noise and QE operations are symmetric was studied.
-Here, we consider the relative frequency of noise and QE operations, $q$, is no longer constrained to be $0.5$.
-The input state is entangled with a reference system $R$, and the entire quantum circuit can be viewed as a quantum channel from $R$ to $SA$.
+In the following, we denote the system qubits undergoing evolution as $S$, the environment qubits as $E$, and the ancilla qubits as $A$.
+The input state is entangled with a reference system $R$, and the entire quantum circuit can be viewed as a quantum channel from $R$ to $SA$, where $SA\equiv S\cup A$.
 Without loss of generality, we consider the scenario where each system qubit is entangled in a Bell pair with a corresponding reference qubit.
@@
 $I_{C}(R>SA)=\mathcal{S}(SA)-\mathcal{S}(SRA),$ (1)
-where $\mathcal{S}(X)$ denoted the entanglement entropy of the subsystem $X$. Our primary focus is to investigate whether a phase transition in $I_{C}(R>SA)$ occurs as $q$ is varied.
-Analytical analysis.
-When random unitary gates are drawn from the Haar measure, the coherent information can be mapped onto the free energy difference of a classical statistical mechanics model under different boundary conditions [55, 56, 57, 58].
+where $SRA\equiv S\cup R\cup A$ and $\mathcal{S}(X)$ denoted the entanglement entropy of the subsystem $X$.
+Our primary focus is to investigate whether a phase transition in $I_{C}(R>SA)$ occurs as $q$ is varied.
+Analytical analysis—When random unitary gates are drawn from the Haar measure, the coherent information can be mapped onto the free energy difference of a classical statistical mechanics model under different boundary conditions using the replica trick [56, 57, 58, 59].
 Specifically, we have
@@
 where
-$\mathcal{Z}_{XA}^{(n,k)}$ $=$ $\sum_{m}\text{Tr}\left(\mathbb{C}^{X}\Lambda_{A}^{(n,k)}\right),\quad\mathcal{F}_{XA}^{(n,k)}=-\text{log}\mathcal{Z}_{XA}^{(n,k)},$
-$\Lambda_{A}^{(n,k)}$ $=$ $\text{Tr}_{A\cup E}\left[\left(\left|\Psi_{m}\right\rangle\left\langle\Psi_{m}\right|\right)^{\otimes Q}\mathbb{C}^{A}\right].$ (3)
-Here, the degrees of freedom of this statistical mechanics model are spins that take values in the permutation group $\mathbb{S}(Q)$ with $Q=nk+1$, and are denoted as $\left|\sigma\right\rangle$.
+$\mathcal{Z}_{XA}^{(n,k)}$ $=$ $\sum_{m}\text{Tr}\left(\mathbb{C}^{X}\Lambda^{(n,k)}\right),\mathcal{F}_{XA}^{(n,k)}=-\text{log}\mathcal{Z}_{XA}^{(n,k)}.$ (3)
 $X=S,SR$.
-$\mathbb{C}$ corresponds to a particular group member in $\mathbb{S}(Q)$, representing to a particular spin orientation.
-Thus, $\Lambda_{A}^{(n,k)}$ can be identified as the bulk partition function, while $\mathbb{C}^{S}$ and $\mathbb{C}^{SR}$ correspond to two different boundary conditions.
-Notice that $S$ and $R$ represents the top and bottom boundary, respectively.
-This statistical mechanics model is a ferromagnetically coupled spin model with random fields oriented in two different directions on a honeycomb lattice [52].
-The vertical bond contributes a weight of $\text{Wg}^{Q}(\sigma^{-1}\tau)$, where $\text{Wg}^{Q}(g)$ is the Weingarten function, acting as a ferromagnetic coupling that assigns greater weights to configurations where $\sigma=\tau$.
+$\Lambda^{(n,k)}$ can be considered the bulk partition function, where the underlying degrees of freedom are spins that take values in the permutation group $\mathbb{S}(Q)$ with $Q=nk+1$.
+$\mathcal{Z}_{SA}^{(n,k)}$ and $\mathcal{Z}_{SRA}^{(n,k)}$ are interpreted as total partition functions under different boundary conditions.
+We denote $\mathbb{I}$ and $\mathbb{C}$ as the identity element and another specific group element in $\mathbb{S}(Q)$, constructed from swap operations, with its precise definition provided in the Supplemental Material [60].
+Specifically, $\mathbb{C}^{S}$ indicates that the top boundary aligns with $\mathbb{C}$ and the bottom with $\mathbb{I}$, while $\mathbb{C}^{SR}$ means both top and bottom boundaries align with $\mathbb{C}$.
+The statistical mechanics model encoded in $\Lambda^{(n,k)}$ is a ferromagnetically coupled spin model with random fields oriented in two different directions on a honeycomb lattice [53].
+The vertical bonds are determined by the random unitary gates, which act as a ferromagnetic coupling that assigns higher weights to configurations where neighboring spins are aligned in the same direction.
 The non-vertical bonds are determined by measurements, QE operations and noise.
-While their exact contributions depend on the specific QE operations and noise model, QE operations generally behave as external fields applied at random locations, aligning in the direction $\mathbb{I}$, corresponding to the identity element in $\mathbb{S}(Q)$.
-Conversely, noise acts as an external field oriented in the direction $\mathbb{C}$.
-Concrete expressions for the bond weights under various noises are provided in Ref.
-[59].
-The parameter $q$ governs the relative frequency of QE operations and noise, thereby tuning the net field direction.
-For small $q$, the net field aligns with $\mathbb{C}$, whereas for large $q$, it aligns with $\mathbb{I}$.
-When combined with the results in Ref.
-[52], the full phase diagram is depicted in Fig.
+Although their exact contributions depend on the specific QE operations and noise model, noises generally behave as external fields applied at random locations, aligning in the direction $\mathbb{I}$.
+In contrast, QE operation acts as an external field oriented in the direction $\mathbb{C}$.
+The parameter $q$ governs the relative frequency of QE operations and noise, thus tuning the direction of net external field.
+Concrete expressions for the bond weights under various noises are provided in [60].
+For small $q$, bulk spins align with $\mathbb{C}$, leading to a domain wall at the bottom boundary when calculating $\mathcal{F}^{(n,k)}_{SA}$, with no additional energy cost in $\mathcal{F}^{(n,k)}_{SRA}$.
+This results in a positive $I_{C}$ and is extensive in system size.
+Since positive $I_{C}$ indicates that finite amount of quantum information can be successfully transmitted through the channel, which can be recovered via a decoding algorithm, we term this the recoverable phase.
+Conversely, for large $q$, bulk spins align with $\mathbb{I}$.
+In this scenario, a domain wall at the top boundary always exists, while an extra domain wall appears in calculating $\mathcal{F}^{(n,k)}_{SRA}$ at the bottom boundary, rendering $I_{C}$ negative and resulting in an irrecoverable phase.
+This coherent information phase transition, tuned by $q$, exhibits characteristics of a first-order phase transition when the sign of $I_{C}$ is considered as the order parameter.
+This contrasts with the MIPT protected by QE operations, which corresponds to a second-order phase transition between ferromagnetic and paramagnetic phases [53].
+The full phase diagram is depicted in Fig.
 1(b).
-By identifying $p$ as the temperature and considering noise and QE operations as two competing external fields, the phase diagram is reminiscent of that of a simple 2D Ising model.
-Consequently, a first-order phase transition is expected when tuning $q$.
-Furthermore, the coherent information can directly detect this first-order phase transition.
-For small $q$, bulk spins align with $\mathbb{C}$, leading to a domain wall at the bottom boundary when calculating $\mathcal{F}^{(n,k)}_{SA}$, with no additional energy cost in $\mathcal{F}^{(n,k)}_{SAR}$.
-This results in a positive $I_{C}$ and is extensive in system size.
-Conversely, for large $q$, bulk spins align with $\mathbb{I}$.
-In this scenario, a domain wall at the top boundary always exists, while an extra domain wall appears in calculating $\mathcal{F}^{(n,k)}_{SAR}$ at the bottom boundary, rendering $I_{C}$ negative.
-Since positive $I_{C}$ indicates that finite amount of quantum information can be successfully transmitted through the channel, which can be recovered via a decoding algorithm, this first-order phase transition corresponds to a transition from a recoverable to an irrecoverable phase.
+By identifying $p$ as the temperature and considering noise and QE operations as two competing external fields, the phase diagram is reminiscent of that of a 2D Ising model.
+A central contribution of our work is the recognition that the additional symmetry-breaking phase induced by QE operations in the classical statistical mechanics mapping corresponds to the recoverable phase, in which quantum information can be reliably transmitted.
 Figure 2: Numerical simulation for coherent information phase transition.
+$L$ denotes number of qubits.
 (a) Resetting noise.
@@
 The circuit is evolved for $5L$ time steps.
-Numerical results.
-To explicitly demonstrate the existence of the coherent information phase transition, we perform numerical simulations on a large scale.
+Numerical results—To explicitly demonstrate the existence of the coherent information phase transition, we perform large-scale numerical simulations in a (1+1)-d quantum circuit.
 In these simulations, random unitary gates are selected from the Clifford group to enable the use of the stabilizer formalism [61, 62, 63].
 We set $U_{SA}=\text{SWAP}$ and specifically consider three types of noise: resetting, depolarizing, and dephasing.
-Here we set $q_{t}=0.1$, while results for other parameters can be found in Supplemental Material [59].
-We first consider the case with $p=0$, and the results are presented in Fig.
-2(a)-2(c).
-A distinct recoverable phase, characterized by positive coherent information, is identified.
+Here we set the total frequency of noise and QE operations to be $q_{t}=0.1$, while results for other parameters can be found in [60].
+We first consider the case with measurement probability $p=0$, and the results are presented in Fig.
+2(a-c).
+A distinct recoverable phase, characterized by positive coherent information, is observed in all cases, highlighting the effectiveness of QE operations in combating various noises.
 Through data collapse, we determine the critical points and critical exponents to be $q_{c}^{\text{reset}}=0.500(1)$, $q_{c}^{\text{depo}}=0.360(1)$ and $q_{c}^{\text{deph}}=0.503(1)$, with $\nu_{c}^{\text{reset}}=0.90(5)$, $\nu_{c}^{\text{depo}}=0.96(5)$ and $\nu_{c}^{\text{deph}}=0.95(7)$, respectively.
-The recoverable phase region is smallest for depolarizing noise and largest for dephasing noise, consistent with the relative strengths of the corresponding random fields [59].
+The recoverable phase region is the smallest for depolarizing noise and the largest for dephasing noise, consistent with the relative strengths of the corresponding random fields [60].
 When measurements are incorporated, the phase transition persists, as demonstrated in Fig.
@@
 As illustrated in Fig.
-3(a), the convergence of the coherent information slows significantly near $q=0.5$ for resetting noise.
+3(a), the convergence of coherent information slows significantly near $q=0.5$ for resetting noise.
 By defining the convergence time $t_{c}$ as the time when $\left|I_{C}(t)-I_{C}(t_{\text{final}})\right|$ first falls below $0.05$, we plot $t_{c}$ as a function of $q$ for different system sizes in Fig.
@@
 The results indicate that $t_{c}$ diverges with system size as $q$ approaches the critical point, suggesting an infinite convergence time in the thermodynamic limit.
-In contrast, in the absence of unitary gates, while the coherent information still transitions from positive to negative precisely at $q_{c}=0.5$ (corresponding to the scenario where half of the system qubits are discarded), critical slowing down was not observed, as shown in Fig.
+In contrast, in the absence of unitary gates, while coherent information still transitions from positive to negative precisely at $q_{c}=0.5$ (corresponding to the scenario where half of the system qubits are discarded), critical slowing down was not observed, as shown in Fig.
 3(c) and Fig.
@@
 This observation underscores the indispensability of unitary gates for the manifestation of a phase transition.
-Efficient protocol.
-Measuring coherent information for a specific circuit and trajectory requires repeated preparation of the same final state.
-However, in realistic experiments, noise locations are usually uncontrollable, making it difficult to reliably reproduce the same circuit.
+Efficient protocol—Measuring coherent information for a specific circuit and trajectory requires repeated preparation of the same final state.
+However, in realistic experiments, the noise locations are usually uncontrollable, making it difficult to reliably reproduce the same circuit.
 This issue is compounded by the post-selection problem when measurements are involved, due to Born’s rule [68, 69, 55].
-Motivated by previous work on constructing post-selection-free probes that reflect the correlation between a quantum device and classical simulations [69, 53, 70, 71, 72], we propose a resource-efficient protocol to observe the coherent information phase transition.
-In the quantum device, an initial state $\rho$ is prepared and encoded using random unitary gates alone.
-This encoded state then undergoes the circuit, incorporating ancilla qubits during the process.
-The unnormalized final state on $A$ is denoted as $\widetilde{\rho_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\rho_{A}^{m}})=p_{\rho}^{m}$, where $p_{\rho}^{m}$ represents the probability of the particular trajectory $m$.
-For brevity, we omit the explicit dependence on the circuit realization $\mathcal{C}$, which includes choices of random unitary gates and locations of noise, QE operations and measurements.
-Leveraging the knowledge of the measurement outcomes and the circuit realization, we simulate the circuit classically with an encoded version of a distinct input state $\sigma$.
-Assuming the unitary gates are Clifford gates and $\sigma$ is a stabilizer state, the classical simulation becomes computationally efficient.
-We then also obtain $\widetilde{\sigma_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\sigma_{A}^{m}})=p_{\sigma}^{m}$ [73].
+Motivated by previous work on constructing post-selection-free probes that reflect the correlation between a quantum device and classical simulations [70, 54, 71, 72, 73], we propose a resource-efficient protocol to observe the coherent information phase transition, utilizing every run of the quantum circuit.
+It is important to note that our protocol relies on the knowledge of noise locations.
+While controlling where such noise occurs is inherently challenging, identifying its location post-occurrence is considerably more tractable.
+Consider running the same circuit realization $\mathcal{C}$ twice, each with distinct initial states $\rho$ and $\sigma$.
+The circuit realization $\mathcal{C}$ includes choices of random unitary gates and locations of noise, QE operations and measurements.
+The unnormalized final state on ancilla qubits $A$ is denoted as $\widetilde{\rho_{A}^{m}}$ and $\widetilde{\sigma_{A}^{m}}$, where $\mathrm{Tr}(\widetilde{\rho_{A}^{m}})=p_{\rho}^{m}$ and $\mathrm{Tr}(\widetilde{\sigma_{A}^{m}})=p_{\sigma}^{m}$, with $p_{\rho}^{m}$ and $p_{\sigma}^{m}$ representing the probability of the particular trajectory $m$ for each respective run.
+For brevity, we omit the explicit dependence of $\widetilde{\rho_{A}^{m}}$ and $\widetilde{\sigma_{A}^{m}}$ on $\mathcal{C}$.
+We can then define the following quantity:
+$\chi=\mathbb{E}_{\mathcal{C}}\frac{\sum_{m}\text{Tr}\left(\widetilde{\rho_{A}^{m}}\widetilde{\sigma_{A}^{m}}\right)}{\sum_{m}\text{Tr}\left(\widetilde{\sigma_{A}^{m}}^{2}\right)},$ (4)
+where $\mathbb{E}_{\mathcal{C}}$ represents averaging over different circuit realizations.
+It resembles the “quantum” cross entropy introduced in Ref.
+[54], with the notable distinction that ancilla qubits introduced dynamically during the evolution, rather than being present from the beginning.
+Furthermore, while $\chi$ does not reveal MIPT, it is a robust indicator of the coherent information transition.
+Intuitively, this probe reflects the distinguishability of different initial states by leveraging access to both the measurement outcomes and the final state of the ancilla qubits.
+In the recoverable phase, the initial information is preserved, and therefore we expect $\chi=0$, corresponding to perfect distinguishability.
+In contrast, in the irrecoverable phase, the initial information is erased, causing $\chi$ to approach 1.
+A rigorous analytical analysis demonstrating the sharp transition of $\chi$ at the phase transition is provided in [60].
 Figure 4: Efficient protocol.
 (a) Schematic workflow of the protocol.
-An important additional ingredient, compared to the approach in [53], is the incorporation of projective measurements on the ancilla qubits.
+An important additional ingredient, compared to the approach in Ref.
+[54], is the incorporation of projective measurements on the ancilla qubits.
 (b) Take depolarizing noise as an example.
@@
 (c) Analogous to (b), with measurement probability $p=0.1$.
-The quantity that serves as a proxy for the phase transition is constructed as follows:
-$\chi$ $=$ $\mathbb{E}_{\mathcal{C}}\frac{\sum_{m}\text{Tr}\left(\widetilde{\rho_{A}^{m}}\widetilde{\sigma_{A}^{m}}\right)}{\sum_{m}\text{Tr}\left(\widetilde{\sigma_{A}^{m}}^{2}\right)}$ (4)
-$=$ $\mathbb{E}_{\mathcal{C}}\sum_{m}p_{\rho}^{m}\frac{p_{\sigma}^{m}}{2^{-N_{\text{rand}}^{\sigma}}}\text{Tr}\left(\rho_{A}^{m}\prod_{k=1}^{l_{\sigma}}\frac{1+g_{k}}{2}\right).$
-Here, $\mathbb{E}_{\mathcal{C}}$ represents averaging over different circuit realizations, $N_{\text{rand}}^{\sigma}$ is the number of measurements with random outcomes during the classical simulation, $l_{\sigma}$ represents the number of generators for the normalized stabilizer state $\sigma_{A}^{m}$ and $g$ are the corresponding generators.
-The middle term $p_{\sigma}^{m}/2^{-N_{\text{rand}}^{\sigma}}$ is either 00 or $1$ due to the property of stabilizer formalism, while the final term corresponds to the probability of measuring all the generators with outcomes equal to $1$.
-For each execution of the quantum circuit, we first compute $p_{\sigma}^{m}/2^{-N_{\text{rand}}^{\sigma}}$ by classical simulation.
-If the result is $1$, we proceed to perform projective measurements on the ancilla qubits according to the stabilizer generators.
+A key advantage of $\chi$ is that its measurement allows us to fully utilize each circuit realization and trajectory, circumventing the need for experimental control over noise locations or post-selection on measurement outcomes.
+The central idea is that the run with initial state $\sigma$ can be classically simulated by selecting it as a stabilizer state.
+When the unitary gates are Clifford gates, $\widetilde{\sigma_{A}^{m}}$ remains a stabilizer state, leading to the following expression for $\chi$:
+$\chi=\mathbb{E}_{\mathcal{C}}\sum_{m}p_{\rho}^{m}\frac{p_{\sigma}^{m}}{2^{-N_{\text{rand}}^{\sigma}}}\text{Tr}\left(\rho_{A}^{m}\prod_{k=1}^{l_{\sigma}}\frac{1+g_{k}}{2}\right).$ (5)
+Here, $N_{\text{rand}}^{\sigma}$ is the number of measurements with random outcomes during the classical simulation, $l_{\sigma}$ represents the number of generators for the normalized stabilizer state $\sigma_{A}^{m}$ and $g$ are the corresponding generators.
+The middle term $p_{\sigma}^{m}/2^{-N_{\text{rand}}^{\sigma}}$ is either 0 or $1$ due to the property of stabilizer formalism, while the final term corresponds to the probability of measuring all the generators with outcomes equal to $1$.
+Thus, the overall process for measuring $\chi$ is outlined in Fig.
+4(a) and is detailed as follows.
+For each execution of the quantum circuit on a quantum device, we first conduct the classical simulation of the same circuit with initial state $\sigma$, leveraging the knowledge of the measurement outcomes and the circuit realization.
+We then compute $p_{\sigma}^{m}/2^{-N_{\text{rand}}^{\sigma}}$ via classical simulation.
+If the result is $1$, we proceed to perform projective measurements on the ancilla qubits in the quantum device according to the stabilizer generators.
 If all measurement outcomes are $1$, we denote this as a successful event.
 The quantity $\chi$ is then estimated as the proportion of successful events among all circuit executions.
-The whole procedure is outlined in Fig.
-4(a).
-Remarkably, this approach allows us to fully leverage each circuit realization and trajectory, circumventing the need for experimental control over noise locations or post-selection on measurement outcomes.
-The proxy $\chi$ resembles the “quantum” cross entropy introduced in Ref.
-[53], with the notable distinction that ancilla qubits are introduced dynamically in this context.
-Furthermore, while this proxy does not reveal MIPT, it proves to be a robust indicator of the coherent information transition.
-As demonstrated in Supplemental Material [59], $\chi$, akin to the coherent information, is directly related to the free energy difference between distinct boundary conditions imposed on the bottom boundary.
-This dependency drives a sharp transition in $\chi$ from 00 to $1$.
-Physically, this probe reflects the distinguishability of different initial states by leveraging access to both the measurement outcomes and the final state of the ancilla qubits.
-In the recoverable phase, the initial information is preserved, and we therefore expect $\chi=0$, corresponding to perfect distinguishability.
-Conversely, in the irrecoverable phase, the initial information becomes erased, causing $\chi$ to approach 1.
 To demonstrate the efficacy of this probe, we present numerical results in Fig.
@@
 4(c) for the case of depolarizing noise with $p=0$ and $p=0.1$, respectively.
+Here, $\rho$ is chosen as a stabilizer state for simulating the quantum device, but it could in principle be any state, rendering the evolution intractable classically [73].
 The critical points are determined to be $q_{c}=0.362(2)$ and $q_{c}=0.378(2)$, which are in agreement with the exact phase transition points.
-Discussions.
-The coherent information is related to the single-shot quantum channel capacity by the relation $\mathcal{Q}_{1}=\max_{\rho_{SR}}I_{C}(R>SA)$ [9, 2].
+Discussions—The coherent information is related to the single-shot quantum channel capacity by the relation $\mathcal{Q}_{1}=\max_{\rho_{SR}}I_{C}(R>SA)$ [8, 2].
 Here we consider $\rho_{SR}$ to be a direct product of Bell pairs between $S$ and $R$.
-Through analytical mapping, it is evident that this choice maximizes the coherent information within the recoverable phase, minimizing the free energy $\mathcal{F}^{(n,k)}_{SAR}$.
+Through analytical mapping, it is evident that this choice maximizes coherent information within the recoverable phase, minimizing the free energy $\mathcal{F}^{(n,k)}_{SAR}$.
 However, in the irrecoverable phase, selecting $\rho_{SR}$ as a trivial product state would yield zero coherent information, thereby maximizing the coherent information.
 This results in a phase transition where the single-shot quantum channel capacity shifts from a finite positive value to zero.
-Furthermore, the true quantum channel capacity $\mathcal{Q}$, which allows for multiple uses of the channel, becomes additive and equal to $\mathcal{Q}_{1}$ only if the channel is degradable [74, 75, 76, 77, 78, 79, 80].
-Although measurement and QE operations are all degradable, general noises often lacks this property, making the calculation of $\mathcal{Q}$ more challenging.
+Generalizing to the true quantum channel capacity $\mathcal{Q}$, which allows for multiple uses of the channel, general noises are non-degradable, making $\mathcal{Q}$ non-additive and complicating the calculation [74, 75, 76, 77, 78, 79, 80].
 Thus, it is intriguing to explore whether a phase transition in $\mathcal{Q}$ can be identified within this framework.
-[…]
+A key aspect of the resource-efficient protocol for detecting the coherent information phase transition is the requirement for knowledge of noise locations.
+In a practical experimental setting, it has been demonstrated that various types of noise can be effectively converted into erasure errors, which can be located by verifying whether the qubit remains in the computational subspace [81, 82, 83, 84, 85, 86, 87, 88, 89].
+Upon detection of an erasure error, replacing the qubit with a maximally mixed state corresponds to a depolarizing noise, while replacing it with $\left|0\right$> corresponds to a resetting noise.
+It is therefore of interest to investigate whether imperfect detection of erasure errors would compromise the protocol’s success.
+We leave this to future work.
 Acknowledgements.
 Acknowledgment.
-We are deeply indebted to Xiao-Liang Qi for many valuable discussions.
+We thank Xiao-Liang Qi and Chong Wang for valuable discussions.
 This work is supported by the Natural Science Foundation of China through Grants No.
@@
 24LZ1400100 and No.
-2019SHZDZX01.
+2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission.